After a recent election, there were 2 more Republicansa governors than Democratic governors in the United States. If there are 50 states governors total, how many Democrats and how many Republicans held governor's offices after this election?
Answer by checkley77(12844)

R+D=50
R=D+2
D+2+D=50
2D=48
D=48/2
D=24 NUMBER OF DEMOCRAT GOVERNORS.
R=24+2=26 NUMBER OF REPUBLICAN GOVERNORS.
PROOF:
24+26=50
50=50